Financial data for Joel de Paris, Inc., for last year follow:

Joel de Paris, Inc.
Balance Sheet
Beginning
Balance
Ending
Balance
Assets
Cash $ 134,000 $ 134,000
Accounts receivable 342,000 489,000
Inventory 576,000 472,000
Plant and equipment, net 808,000 805,000
Investment in Buisson, S.A. 404,000 429,000
Land (undeveloped) 252,000 248,000
Total assets $ 2,516,000 $ 2,577,000
Liabilities and Stockholders' Equity
Accounts payable $ 373,000 $ 348,000
Long-term debt 998,000 998,000
Stockholders' equity 1,145,000 1,231,000
Total liabilities and stockholders' equity $ 2,516,000 $ 2,577,000


Joel de Paris, Inc.
Income Statement
Sales $ 4,324,000
Operating expenses 3,632,160
Net operating income 691,840
Interest and taxes:
Interest expense $ 126,000
Tax expense 195,000 321,000
Net income $ 370,840


The company paid dividends of $284,840 last year. The “Investment in Buisson, S.A.,” on the balance sheet represents an investment in the stock of another company. The company's minimum required rate of return of 15%.

Required:

1. Compute the company's average operating assets for last year.

2. Compute the company’s margin, turnover, and return on investment (ROI) for last year. (Round "Margin", "Turnover" and "ROI" to 2 decimal places.)

3. What was the company’s residual income last year?
Answer #1

Average operating assets

= [(2,516,000-252,000-404,000)+(2,577,000-248,000-429,000)]/2

= 1,880,000

.

Margin = Net operating income/Sales = 691,840/4,324,000

= 16%

.

Turnover = Sales/Average operating assets = 4,324,000/1,880,000

= 2.3

.

ROI = Margin * Turnover = 16*2.3

= 36.8%

.

Residual income = Net operating income - required return

= 691,840 - (1,880,000*15%)

= 409,840